Find the radian measure of the largest angle in the triangle with side lengths of 77​, 88​, and 1313.
Answer by Alan3354(69443)

Find the radian measure of the largest angle in the triangle with side lengths of 77​, 88​, and 1313.
---------
I'll assume you mean 7, 8 & 13
===
The largest angle is opposite the longest side.
Use the Cosine Law
13^2 = 7^2 + 8^2 - 2*7*8*cos(C)
cos(C) = 56/112 = 0.5
C = pi/3